# Improved Belief Propagation Decoding algorithms for surface codes

### v2

@@ III.3 Complexity
Our proposed algorithms involve the same message passing between variable nodes and check nodes as traditional BP. The main difference lies in the additional update steps, which in Momentum-BP and AdaGrad-BP consist of recalculating the a posteriori probabilities, and in EWAInit-BP involves recalculating the a priori probabilities. These additional steps only involve simple arithmetic operations and can be executed locally at each node, without requiring extra iterations, loops, or global communication, thus not introducing any significant overhead. Therefore, the overall time complexity of the proposed algorithms remains $O(j\tau N)$, where $j$ is the average degree of the Tanner graph, and $\tau$ is the number of iterations.
@@ IV Simulation Results
Our simulations employ a depolarizing channel, where each qubit has an equal probability $p/3$ of experiencing Pauli $X$, $Y$, or $Z$ errors, and a probability of $1-p$ of no error occurring. Additionally, for ease of comparison, we adopt the simple noise model which does not consider measurement errors or error propagation in quantum circuits.
@@ V Conclusion
In this article, we proposed three improved BP algorithms which are Momentum-BP, AdaGrad-BP and EWAInit-BP.
Since the a posteriori probability update of BP can be interpreted as the gradient optimization of an energy function, we borrowed optimization strategies from gradient descent in machine learning, proposing Momentum-BP and AdaGrad-BP. These mechanisms aim to smooth the update process and reduce the amplitude of oscillations, respectively. Experiments on trapping sets intuitively demonstrated the effectiveness of these algorithms.
From the theoretical derivation of the message update rules of BP algorithm, we realized the importance of a priori probabilities and transformed the message update smoothing into initial values adapting, proposing EWAInit-BP. This algorithm exhibits aggressive exploration and outperforms other algorithms without OSD post-processing on the toric code, planar surface code, and XZZX surface code. Specifically, EWAInit-BP can demonstrate error correction capabilities even under parallel scheduling, with logical error rates decreasing as code length increases. Notably, all our proposed methods do not introduce any additional factors to the time complexity of BP. Thus, EWAInit-BP can be implemented with $O(1)$ complexity in fully parallel while maintaining decoding accuracy.
However, the aggressive exploration and convergence capabilities of EWAInit increase the probability of generating logical operators after error recovery. Therefore, directly applying this method to other error-correcting codes presents certain challenges. A potential research direction is to identify situations where logical operators may occur. Besides, our simulations only considered Pauli errors on qubits and did not account for measurement errors, error propagation in circuits, etc. Future work needs to adapt BP to more challenging noise models, potentially requiring specialized optimizations.
@@ Acknowledgements.
We would like to express our sincere gratitude to Zhengzhong Yi for his in-depth discussions on the principles of BP algorithms and meticulous feedback on the manuscript. We are also thankful to Zhipeng Liang for his assistance with the experimental code framework. Finally, we extend our appreciation to Prof. Xuan Wang for his insightful discussions and guidance on the gradient approach.

### v3

@@ III.3 Complexity
Our proposed algorithms involve the same message passing between variable nodes and check nodes as in traditional BP. The main difference lies in the additional update steps, which in Momentum-BP and AdaGrad-BP consist of recalculating the a posteriori probabilities, and in EWAInit-BP involves recalculating the a priori probabilities. These additional steps only involve simple arithmetic operations and can be executed locally at each node, without requiring extra iterations, loops, or global communication, thus not introducing any significant overhead. Therefore, the overall time complexity of the proposed algorithms remains $O(j\tau N)$, where $j$ is the average degree of the Tanner graph, and $\tau$ is the number of iterations.
@@ IV Simulation Results
Without further specification, our simulations employ a depolarizing channel, where each qubit has an equal probability $p/3$ of experiencing Pauli $X$, $Y$, or $Z$ errors, and a probability of $1-p$ of no error occurring. Additionally, for ease of comparison, we adopt the simple noise model which does not consider measurement errors or error propagation in quantum circuits.
@@ V Conclusion
In this article, we proposed three improved BP algorithms: Momentum-BP, AdaGrad-BP and EWAInit-BP.
Since the a posteriori probability update of BP can be interpreted as the gradient optimization of an energy function, we adopted optimization strategies from gradient descent in machine learning, resulting in Momentum-BP and AdaGrad-BP. These mechanisms aim to smooth the update process and reduce the amplitude of oscillations. Simulations on trapping sets intuitively demonstrated the effectiveness of these algorithms.
From the theoretical derivation of the message update rules of BP algorithm, we recognized the importance of a priori probabilities and transformed the message update smoothing into initial values adapting, proposing EWAInit-BP. This algorithm exhibits aggressive exploration and outperforms other algorithms without OSD post-processing on the toric code, planar surface code, and XZZX surface code. Specifically, EWAInit-BP shows high decoding accuracy even under parallel scheduling, with logical error rates decreasing as code length increases. Notably, all our proposed methods do not introduce any additional factors to the time complexity of BP. Thus, EWAInit-BP can be implemented with $O(1)$ complexity in fully parallelism while maintaining decoding accuracy.
However, the aggressive exploration and convergence capabilities of EWAInit-BP increase the probability of generating logical operators after error recovery. Therefore, directly applying this method to other error-correcting codes presents certain challenges. A potential research direction is to identify situations where logical operators may occur. Besides, our simulations only considered Pauli errors on qubits and did not account for measurement errors, error propagation in circuits, etc. Future work needs to adapt BP to more challenging noise models, potentially requiring specialized optimizations.
@@ Acknowledgements.
This work was supported by the Colleges and Universities Stable Support Project of Shenzhen, China (No.GXWD20220817164856008) and Harbin Institute of Technology, Shenzhen - SpinQ quantum information Joint Research Center Project (No.HITSZ20230111). We would like to express our sincere gratitude to Zhengzhong Yi for his in-depth discussions on the principles of BP algorithms and meticulous feedback on the manuscript. We are also thankful to Zhipeng Liang for his assistance with the experimental code framework. Finally, we extend our appreciation to Prof. Xuan Wang for his insightful discussions and guidance on the gradient approach.